OpenStudy (caitlinnr14):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (texaschic101):

use point slope formula because you have the slope and the point y - y1 = m(x - x1) m = -6 (-2,-3) x1 = -2, y1 = -3 now lets sun our info in y - (-3) = -6(x - (-2) can you take it from here ?

OpenStudy (caitlinnr14):

im lost @texaschic101

OpenStudy (texaschic101):

y - (-3) = -6(x - (-2) y + 3 = -6(x + 2) y + 3 = -6x - 12 y = -6x - 12 - 3 y = -6x - 15 Do you understand ?